# DEM: Distribution Edited Model for Training with Mixed Data Distributions

### v1

@@ Abstract
Training with mixed data distributions is a common and important part of creating multi-task and instruction-following models. The diversity of the data distributions and cost of joint training makes the optimization procedure extremely challenging. Data mixing methods partially address this problem, albeit having a sub-optimal performance across data sources and require multiple expensive training runs. In this paper, we propose a simple and efficient alternative for better optimization of the data sources by combining models individually trained on each data source with the base model using basic element-wise vector operations. The resulting model, namely Distribution Edited Model (DEM), is $\times1$ cheaper than standard data mixing and outperforms strong baselines on a variety of benchmarks, yielding upto 6.2% improvement on MMLU, 11.5% on BBH, 16.1% on DROP, and 9.3% on HELM with models of size 3B to 13B. Notably, DEM does not require full re-training when modifying a single data-source, thus making it very flexible and scalable for training with diverse data sources.
@@ 1 Introduction
Large Language Models (LLM) go through an extensive pretraining on billions or trillions of tokens (Brown et al., 2020; Zhang et al., 2022; Raffel et al., 2020; Touvron et al., 2023a, b; Geng and Liu, 2023), but they typically require supervised fine-tuning on diverse instruction-following datasets for properly following human instructions Ouyang et al. (2022); Chung et al. (2022); Sanh et al. (2022); Iyer et al. (2022). Supervised training is crucial for ensuring that generated outputs meet user expectations and perform well on downstream tasks Radford et al. (2019); Gururangan et al. (2020).
The datasets for supervised training are often of different sizes and follow different distributions. Recent state-of-the-art fine-tuning approaches Chung et al. (2022); Iyer et al. (2022) demonstrate that training on multiple data distributions requires careful tuning of the mixing weights for each data source to capture the combined distribution and improve downstream task performance. Tuning these weights in a data-mixing approach is a computationally expensive process. Although, there are techniques to speed-up the search (Xie et al., 2023; Albalak et al., 2023), the process remains time-consuming. Moreover, when one or more new datasets are introduced, the weights for each dataset need to be re-tuned. This requirement makes the data-mixing approach inflexible and hard to maintain in a production environment.
To address these challenges when fine-tuning a LLM on a set of diverse data distributions, we propose a simple and efficient approach that combines individually trained versions of the base model using element-wise vector operations. Our method focuses on the challenging setting of combining diverse data distributions that correspond to multiple tasks from different domains such as math reasoning, conversations and coding. In particular, our goal is to better capture a diverse data distributions as opposed to editing the model on a single downstream task Ilharco et al. (2022a); Schumann et al. (2023). Hence, we call resulting model Distribution Edited Model (DEM, shown in Figure 1). Our experiments on a variety of downstream tasks show that DEMs are an effective, highly capable and low cost alternatives to the models trained using data mixing methods.
@@
  • We show that DEM reduces the training cost by $\times1$ while improving the model performance. Compared to standard data mixing approaches, DEM yields up to 6.2% improvement on MMLU, 11.5% on BBH, 16.1% on DROP, and 9.3% on HELM with 3B, 7B, and 13B models.
@@ 2 Related Work
Instruction-based multi-task fine-tuning of language models has been previously shown to improve both zero and few-shot performance on unseen tasks  (Wei et al., 2022a; Sanh et al., 2022). Instruction-tuning data can be sourced from diverse task categories (such as math, reasoning, dialog etc), and the model performance is often sensitive to the data-mixing strategy. For example, both  (Chung et al., 2022) and  (Iyer et al., 2022) carefully tune the data-mixing weights for various training data sources.
@@
Recently, model weight interpolation and task arithmetic techniques have been shown to improve the performance of pre-trained models on: single-task Izmailov et al. (2018); Matena and Raffel (2022); Yüce et al. (2022); Wortsman et al. (2022b) and multi-task Ilharco et al. (2022b, a); Li et al. (2022); Wortsman et al. (2022a); Yadav et al. (2023); Daheim et al. (2023), out-of-domain generalization Cha et al. (2021); Arpit et al. (2022); Rame et al. (2022); Jin et al. (2023); Ramé et al. (2023), and federated learning McMahan et al. (2017); Li et al. (2020).
@@ 5.1 Dataset
Here, we list the fine-tuning datasets, we use to enhance instruction following capability of our base pre-trained LLM. Previous work has shown that they improve the instruction following capabilities of the models (Chung et al., 2022; Iyer et al., 2022; Gupta et al., 2022; Amini et al., 2019; Sanh et al., 2022).
@@ 5.3 Training
We fine-tune the OpenLLaMA model on all instruction following datasets (Section 5.1), both separately and jointly. We use AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.95$, weight decay of 0.05, gradient clipping of 1 and a constant learning rate of 2e-5 with a 2000 step warmup. We also adjust batch size for different datasets based on the validation loss (see Appendix A for details). We use a greedy sample packing approach to fit multiple training samples into a single batch sample efficiently, padding to the max sequence length without overflowing into the next sample of a batch. To select the optimal mixing weights for DEM (Eq 2), we perform a grid search over $\omega_{i}$ values. For each coefficient combination we evaluate the validation losses on the five datasets (Section 5.1), and select the model that minimizes their average (see Appendix B for details).
@@ 5.4 Evaluation Framework
We evaluate the instruction following capability of the models using two publicly available benchmarks, namely InstructEval (Chia et al., 2023) and HELM (Liang et al., 2023). To have a holistic evaluation, we choose a diverse set of held-out datasets: (i) from InstructEval – MMLU, Big-Bench Hard and DROP, and (ii) from HELM – twenty sets from six diverse task-groups – Classification, ClosedbookQA, OpenbookQA, Math, Reasoning and Conversational (see Table 3). We perform 5-shot evaluation on MMLU and HELM, and 3-shot evaluation on BBH and DROP, inline with the standardized setup and previous work.
@@ 5.5 Baseline Models
The pre-trained OpenLLaMA serves as the non instruction-tuned baseline for evaluation. Our primary instruction-tuned baseline is data mixing – the model is fine-tuned using a weighted mixture of 5 diverse datasets as described in 5.1 following Chung et al. (2022); Iyer et al. (2022) which has been shown to produce SOTA performance with large scale diverse datasets. This model requires finding the optimal weights corresponding to each training dataset such that the validation loss reaches optimal value for each dataset at similar number of training steps. We experimented with several combinations of weights and chose the one that leads to the smallest validation loss (see Appendix C for details). Additionally, we create a simpler baseline where we concatenate all 5 training datasets and the samples are shuffled randomly during training. This technique is more cost-effective than the standard data mixing approach, as it does not require any weight optimization.
@@ 6.1 Downstream Task Performance
Models MMLU BBH DROP
Open LLaMA 40.31 32.84 24.38
LLaMA Touvron et al. (2023a) 35.10 30.30 -
LLaMA2 Touvron et al. (2023b) 45.30 32.60 -
Flan PaLM (8B) Chung et al. (2022) 49.3 36.4 -
OPT-IML (30B) Iyer et al. (2022) 43.2 30.9 -
OPT-IML (175B) Iyer et al. (2022) 47.1 35.7 -
CoT 41.67 33.98 24.20
Math QA 39.71 32.70 24.31
P3 35.69 14.00 23.29
InstDial 39.31 23.09 21.81
SNI 46.55 35.88 34.53
Data Mixing 47.77 36.38 32.71
Concatenated Datasets 43.43 21.34 23.21
DEM - Interpolation (Ours) 50.14 40.11 36.31
DEM - Distribution Vector (Ours) 50.74 40.56 37.96
Table 2: Downstream task performance of models trained on different instruction following datasets (Section 5.1). We compare it with different pretrained and fine-tuned baselines (Section 5.5) and our proposed approach in Section 4. The models are of size 7B, unless specified. The performance numbers for models with citation are taken from the corresponding paper, rest are evaluated using InstructEval.
@@
We present the performance of fine tuned models on each dataset separately and observe that the performance degrades compared to Open-LLaMA model for P3, InstDial and MathQA. On the other hand, we observe significant improvement with CoT and SNI datasets in all three task families. We compare these models with data mixing baseline and note that it performs significantly better than the pretrained OpenLLaMA, while the improvemnet compared to the best single dataset fine-tuned model (i.e. SNI) is much smaller, even worse for DROP. The concatenated datasets baseline performs significantly worse than data mixing method, only improving for MMLU compared to OpenLLaMA and significantly worse than SNI fine-tuned model. This highlights the importance of choosing the optimal weights for data mixing and training a strong baseline.
@@
Table 3: Summary results of the HELM evaluations on held-out scenarios, grouped by task-category.
@@ 6.2 Effect of Model Size Scaling
# Params Models MMLU BBH DROP
3B Data Mixing 41.08 31.36 25.98
DEM 43.67 34.14 28.89
7B Data Mixing 47.77 36.38 32.71
DEM 50.74 40.56 37.96
13B Data Mixing 52.7 40.48 43.15
DEM 54.53 42.65 46.59
@@
We evaluate the performance of the proposed DEM approach with increasing model sizes using Open-LLaMA 3B, 7B, and 13B models, quantifying the impact with both smaller and larger models. We trained the baseline Data Mixing model using the method discussed in Section 5.5. On the other hand, we fine-tuned the models on each dataset separately and combined them using Eq 2, similar to the 7B model as discussed in Section 6.1. The results are presented in Table 4. We observe that the model performance increases as we scale up the model size from 3B to 13B for both Data Mixing and DEM. Additionally, DEM yields performance improvement for each model size, showing the effectiveness and generalizability of the proposed approach with model size.
@@ 6.3 Impact of Different Training Datasets
In this section, we analyze the impact of each training dataset included in the mixture on the downstream task performance. For this, we progressively add the data distribution vector corresponding to each dataset to the base model (following Eq. 2) and evaluate the resulting model. We use $\omega_{i}=0.25$ for all datasets to keep the setup simple. The performance of the resulting models are presented in Table 5. We observe that these data sources yield different levels of performance gains, as expected. This can be due to the various levels of mismatch between the train and test distribution. We observe that combining the pretrained model with single-task distribution vectors (e.g Math QA) or smaller mix of tasks (e.g. CoT) leads to smaller improvement whereas large scale multi-task distribution vectors (e.g. P3 and SNI) yields a much larger performance gain, in comparison. It can also be due to the large diversity of tasks and samples in P3 and SNI. InstructDial is an exception, which can be due to the conversational nature of this dataset, making it very different from the evaluation tasks.
Training Dataset MMLU BBH DROP
Open LLaMA 40.31 32.84 24.38
+ CoT 41.30 33.68 25.46
+ MathQA 41.67 33.73 26.05
+ P3 47.12 36.58 30.82
+ InstDial 47.44 38.20 31.15
+ SNI 50.74 40.56 37.96
@@ 6.5 Compute Cost Comparison
We use Eq. 4 to compare the real compute cost of the proposed approach with the baseline data mixing method on Nvidia A100 machines (with 8 gpus per node). In Table 8, we present the gpu-hours used by different training runs, as well as the validation runs needed for finding optimal mixing weight $\omega_{i}$ in Eq 2. In each case, we did early stopping to obtain the best validation loss, which results in varying number of training steps (‘# steps’ in Table 8). As discussed in Appendix C, we use a combination of 10 weights to get the best model for DEM, which costs 23 gpu-hours. The total cost (training+validation) for DEM is 966 gpu-hours.
For the baseline data mixing, we trained 50 models with different weight combination (the exact weight selection process is described in Appendix C). Each run costs an average of 233 gpu-hours, resulting in a total cost of 11650 gpu-hours. This is more than 11 times the total cost of DEM.
@@
Table 8: Training cost (in gpu-hours) on different instruction following datasets computed using Eq 4. Note that the number of steps is not equal to the number of examples.
@@ 7 Conclusions and Future Work
  • Architecture type. The proposed approach makes no specific assumptions regarding the architecture and should be, in principle, applicable to any architecture variant including Mixture-of-Expert models Fedus et al. (2022); Xue et al. (2024); Jiang et al. (2024). Due to budget constraints, the evaluation of different architecture types was not included in the experiment plan. Therefore, the compatibility of DEM with different architecture types remains to be evaluated.

### v2

@@ Abstract
Training with mixed data distributions is a common and important part of creating multi-task and instruction-following models. The diversity of the data distributions and cost of joint training makes the optimization procedure extremely challenging. Data mixing methods partially address this problem, albeit having a sub-optimal performance across data sources and require multiple expensive training runs. In this paper, we propose a simple and efficient alternative for better optimization of the data sources by combining models individually trained on each data source with the base model using basic element-wise vector operations. The resulting model, namely Distribution Edited Model (DEM), is $\times1$ cheaper than standard data mixing and outperforms strong baselines on a variety of benchmarks, yielding upto 6.2% improvement on MMLU, 11.5% on BBH, 16.1% on DROP, 6% on MathQA, and 9.3% on HELM with models of size 3B to 13B. Notably, DEM does not require full re-training when modifying a single data-source, thus making it very flexible and scalable for training with diverse data sources.
@@ 1 Introduction
Large Language Models (LLM) go through an extensive pretraining on billions or trillions of tokens (Brown et al., 2020; Zhang et al., 2022; Raffel et al., 2020; Touvron et al., 2023a, b; Geng and Liu, 2023), but they typically require supervised fine-tuning on diverse instruction-following datasets for properly following human instructions Ouyang et al. (2022); Sanh et al. (2022); Iyer et al. (2022); Chung et al. (2024). Supervised training is crucial for ensuring that generated outputs meet user expectations and perform well on downstream tasks Radford et al. (2019); Gururangan et al. (2020).
The datasets for supervised training are often of different sizes and follow different distributions. Recent state-of-the-art fine-tuning approaches Iyer et al. (2022); Chung et al. (2024) demonstrate that training on multiple data distributions requires careful tuning of the mixing weights for each data source to capture the combined distribution and improve downstream task performance. Tuning these weights in a data-mixing approach is a computationally expensive process. Although, there are techniques to speed-up the search (Xie et al., 2023; Albalak et al., 2023), the process remains time-consuming. Moreover, when one or more new datasets are introduced, the weights for each dataset need to be re-tuned. This requirement makes the data-mixing approach inflexible and hard to maintain in a production environment.
To address these challenges when fine-tuning an LLM on a set of diverse data distributions, we propose a simple and efficient approach that combines individually trained versions of the base model using element-wise vector operations. Our method focuses on the challenging setting of combining diverse data distributions that correspond to multiple tasks from different domains such as math, reasoning, conversations and coding. In particular, our goal is to better capture a diverse data distributions as opposed to editing the model on a single downstream task Ilharco et al. (2022a); Schumann et al. (2024). Hence, we call resulting model Distribution Edited Model (DEM, shown in Figure 1). Our experiments on a variety of downstream tasks show that DEM is an effective, highly capable and low cost alternative to the models trained using data mixing methods.
@@
  • We show that DEM reduces the training cost by $\times1$ while improving the model performance. Compared to standard data mixing approaches, DEM yields up to 6.2% improvement on MMLU, 11.5% on BBH, 16.1% on DROP, 6% on MathQA and 9.3% on HELM with 3B, 7B, and 13B models.
@@ 2 Related Work
Instruction-based multi-task fine-tuning of language models has been previously shown to improve both zero and few-shot performance on unseen tasks  (Wei et al., 2022a; Sanh et al., 2022). Instruction-tuning data can be sourced from diverse task categories (such as math, reasoning, dialog etc), and the model performance is often sensitive to the data-mixing strategy. For example, both  (Chung et al., 2024) and  (Iyer et al., 2022) carefully tune the data-mixing weights for various training data sources.
@@
Recently, model weight interpolation and task arithmetic techniques have been shown to improve the performance of pre-trained models on: single-task Izmailov et al. (2018); Matena and Raffel (2022); Yüce et al. (2022); Wortsman et al. (2022b) and multi-task Ilharco et al. (2022b, a); Li et al. (2022); Wortsman et al. (2022a); Yadav et al. (2023); Daheim et al. (2024), out-of-domain generalization Arpit et al. (2022); Rame et al. (2022); Jin et al. (2023); Ramé et al. (2023); Cha et al. (2024), and federated learning McMahan et al. (2017); Li et al. (2020).
@@ 5.1 Dataset
Here, we list the fine-tuning datasets, we use to enhance instruction following capability of our base pre-trained LLM. Previous work has shown that they improve the instruction following capabilities of the models (Chung et al., 2024; Iyer et al., 2022; Gupta et al., 2022; Amini et al., 2019; Sanh et al., 2022).
@@ 5.3 Training
We fine-tune the OpenLLaMA model on all instruction following datasets (Section 5.1), both separately and jointly. We use AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.95$, weight decay of 0.05, gradient clipping of 1 and a constant learning rate of 2e-5 with a 2000 step warmup. We also adjust batch size for different datasets based on the validation loss (see Appendix A for details). We use a greedy sample packing approach to fit multiple training samples into a single batch sample efficiently, padding to the max sequence length without overflowing into the next sample of a batch. To select the optimal mixing weights for DEM (Eq 2), we perform a grid search over $\omega_{i}$ values. For each coefficient combination we evaluate the validation losses on the five datasets (Section 5.1), and select the model that minimizes their average (see Section C for details). We use an equal weight of $\omega_{i}=0.25$ for all datasets in our experiments.
@@ 5.4 Evaluation Framework
We evaluate the instruction following capability of the models using three publicly available benchmarks, namely InstructEval (Chia et al., 2024), LM-evaluation-harness (Gao et al., 2024) and HELM (Liang et al., 2023). To have a holistic evaluation, we choose a diverse set of held-out datasets: (i) from InstructEval – MMLU, Big-Bench Hard and DROP, (ii) from LM-evaluation-harness – MathQA, and (iii) from HELM – twenty sets from six diverse task-groups – Classification, ClosedbookQA, OpenbookQA, Math, Reasoning and Conversational (see Table 3). We perform 5-shot evaluation on MMLU and HELM, and 3-shot evaluation on BBH and DROP, inline with the standardized setup and previous work.
@@ 5.5 Baseline Models
The pre-trained OpenLLaMA serves as the non instruction-tuned baseline for evaluation. Our primary instruction-tuned baseline is data mixing – the model is fine-tuned using a weighted mixture of 5 diverse datasets as described in 5.1 following Chung et al. (2024); Iyer et al. (2022) which has been shown to produce SOTA performance with large scale diverse datasets. This model requires finding the optimal weights corresponding to each training dataset such that the validation loss reaches optimal value for each dataset at similar number of training steps. We experimented with several combinations of weights and chose the one that leads to the smallest validation loss (see Appendix B for details). Additionally, we create a simpler baseline where we concatenate all 5 training datasets and the samples are shuffled randomly during training. This technique is more cost-effective than the standard data mixing approach, as it does not require any weight optimization.
@@ 6.1 Downstream Task Performance
Models MMLU BBH DROP MathQA
Open LLaMA 40.31 32.84 24.38 27.71
LLaMA Touvron et al. (2023a) 35.10 30.30 - -
LLaMA2 Touvron et al. (2023b) 45.30 32.60 - -
FlanPaLM (8B) Chung et al. (2024) 49.3 36.4 - -
OPT-IML (30B) Iyer et al. (2022) 43.2 30.9 - -
OPT-IML (175B) Iyer et al. (2022) 47.1 35.7 - -
CoT 41.67 33.98 24.20 29.31
Math QA 39.71 32.70 24.31 25.03
P3 35.69 14.00 23.29 27.14
InstDial 39.31 23.09 21.81 26.60
SNI 46.55 35.88 34.53 28.31
Data Mixing 47.77 36.38 32.71 30.35
Concatenated Datasets 43.43 21.34 23.21 27.71
DEM - Interpolation (Ours) 50.14 40.11 36.31 31.22
DEM - Distribution Vector (Ours) 50.74 40.56 37.96 32.16
Table 2: Downstream task performance of models trained on different instruction following datasets (Section 5.1). We compare it with different pretrained and fine-tuned baselines (Section 5.5) and our proposed approach in Section 4. The models are of size 7B, unless specified. The performance numbers for models with citation are taken from the corresponding paper, rest are evaluated using InstructEval and LM-evaluation-harness.
@@
We present the performance of fine tuned models on each dataset separately and observe that the performance degrades compared to Open-LLaMA model for P3, InstDial and MathQA. On the other hand, we observe significant improvement with CoT and SNI datasets in all four task families. We compare these models with data mixing baseline and note that it performs significantly better than the pretrained OpenLLaMA, while the improvement compared to the best single dataset fine-tuned model (i.e. SNI) is much smaller, even worse for DROP. The concatenated datasets baseline performs significantly worse than data mixing method, only improving for MMLU compared to OpenLLaMA and significantly worse than SNI fine-tuned model. This highlights the importance of choosing the optimal weights for data mixing and training a strong baseline.
@@
Table 3: Summary results of the HELM evaluations on held-out scenarios, grouped by task-category for the 7B model. DEM outperforms data mixing approach in five out of six HELM task clusters.
@@ 6.2 Effect of Model Size Scaling
# Params Models MMLU BBH DROP MathQA
3B Data Mixing 41.08 31.36 25.98 28.54
DEM 43.67 34.14 28.89 29.78
7B Data Mixing 47.77 36.38 32.71 30.35
DEM 50.74 40.56 37.96 32.16
13B Data Mixing 52.7 40.48 43.15 30.72
DEM 54.53 42.65 46.59 33.13
@@
We evaluate the performance of the proposed DEM approach with increasing model sizes using Open-LLaMA 3B, 7B, and 13B models, quantifying the impact with both smaller and larger models. We trained the baseline Data Mixing model using the method discussed in Section 5.5. On the other hand, we fine-tuned the models on each dataset separately and combined them using Eq 2, similar to the 7B model as discussed in Section 6.1. We use the same model mixing weight of $\omega_{i}=0.25$ (optimized for 7B model) for models of all sizes and present the results in Table 4. We observe that the model performance increases as we scale up the model size from 3B to 13B for both Data Mixing and DEM. Additionally, DEM yields performance improvement for each model size, showing the effectiveness and generalizability of the proposed approach with model size.
@@ 6.3 Impact of Different Training Datasets
In this section, we analyze the impact of each training dataset included in the mixture on the downstream task performance. For this, we progressively add the data distribution vector corresponding to each dataset to the base model (following Eq. 2) and evaluate the resulting model. We use $\omega_{i}=0.25$ for all datasets to keep the setup simple. The performance of the resulting models are presented in Table 5. We observe that these data sources yield different levels of performance gains, as expected. This can be due to the various levels of mismatch between the train and test distribution. We observe that combining the pretrained model with single-task distribution vectors (e.g Math QA) or smaller mix of tasks (e.g., CoT) leads to smaller improvement whereas large scale multi-task distribution vectors (e.g., P3 and SNI) yields a much larger performance gain, in comparison. It can also be due to the large diversity of tasks and samples in P3 and SNI. InstructDial is an exception, which can be due to the conversational nature of this dataset, making it very different from the evaluation tasks.
Training Dataset MMLU BBH DROP MathQA
Open LLaMA 40.31 32.84 24.38 27.71
+ CoT 41.30 33.68 25.46 28.44
+ MathQA 41.67 33.73 26.05 28.68
+ P3 47.12 36.58 30.82 30.35
+ InstDial 47.44 38.20 31.15 30.65
+ SNI 50.74 40.56 37.96 32.16
@@ 6.5 Compute Cost Comparison
We use Eq. 4 to compare the real compute cost of the proposed DEM approach with the baseline data mixing method for 7B model on Nvidia A100 machines (with 8 gpus per node). Note that, this cost is specific to our setup and it can change depending on the model size, training parallelization scheme and other factors. In Table 8, we present the gpu-hours used by different training runs, as well as the validation runs needed for finding optimal model mixing weight $\omega_{i}$ in Eq 2. In each case, we did early stopping to obtain the best validation loss, which results in varying number of training steps (‘# steps’ in Table 8). As discussed in Appendix B, we use a combination of 10 weights to get the best model for DEM, which costs 23 gpu-hours. The total cost (training+validation) for DEM is 966 gpu-hours.
For the baseline data mixing, we trained 50 models with different weight combination (the exact weight selection process is described in Appendix B). Each run costs an average of 233 gpu-hours, resulting in a total cost of 11650 gpu-hours. This is more than 11 times the total cost of DEM.
@@
Table 8: Training cost (in gpu-hours) of 7B model on different instruction following datasets computed using Eq 4. Note that the number of steps is not equal to the number of examples.
@@ 7 Conclusions and Future Work
Acknowledgments
We thank the anonymous reviewers for their helpful questions and comments, which have helped us improve the quality of the paper. We also want to thank Yang Li for their help in setting up HELM evaluation framework, and Thomas Müller and Lluís Màrquez for helpful discussions.
@@ Acknowledgments
  • Architecture type. The proposed approach makes no specific assumptions regarding the architecture and should be, in principle, applicable to any architecture variant including Mixture-of-Expert models Fedus et al. (2022); Xue et al. (2024); Jiang et al. (2024); Sukhbaatar et al. (2024); Hu et al. (2024). Due to budget constraints, the evaluation of different architecture types was not included in the experiment plan. Therefore, the compatibility of DEM with different architecture types remains to be evaluated.